equation of the pair of lines passing through the orgin whose sum and product of slopes are A.M. and G.M. of 4 and 16 is ?

Grade:9

1 Answers

Harsh Patodia IIT Roorkee
6 years ago
AM of 4,16 is=10
GM of 4,16 is =8
m1+m2=10
m1m2=8 so equation will be (y-m1x)(y-m2x)=0
y2-(m1+m2)xy+m1m2x2=0
y2-10xy+8x2=0